# Channel Estimation for RIS-Assisted mmWave Systems via Diffusion Models

### v2

@@ III-C Network Architecture Design
1) Prepocess: To facilitate network processing, the complex-valued matrix $\mathbf{X}_{t}\in\mathbb{C}^{M\times K}$ is first decomposed into its real part $\Re(\mathbf{X}_{t})$ and imaginary part $\Im(\mathbf{X}_{t})$, and then rearranged into a three-dimension (3D) tensor $\mathbf{X}^{real}_{t}\in\mathbb{R}^{2\times M\times K}$.
[…]
3) Encoder: Motivated by the architectural efficiency and denoising effectiveness of DnCNN [17], we adapt its design principles to channel estimation and introduce a novel BRC block. Specifically, this block consists of Batch normalization (BN) layers, ReLU activation functions, and Convolutional layers to extract multi-level features efficiently. To further enhance the model’s capability, we integrate the self-attention mechanism. This tailored design not only improves denoising performance but also reduces model complexity, making it well-suited for practical wireless communication scenarios.
Temporal information $t$ is embedded into each BRC block via a sinusoidal positional embedding vector $\mathbf{e}\in\mathbb{R}^{N_{t_{embed}}}$, enabling the model to capture temporal dependencies. The embedding vector is then projected into $\mathbf{e}_{project}\in\mathbb{R}^{N_{project}}$ to facilitate subsequent computations by a projection module consisting of a Swish activation function and a linear layer.
@@ IV-A Experiment Setup
We evaluate the proposed method under four scenarios, combining two carrier frequencies ($f_{c}=28$ GHz and $f_{c}=73$ GHz) with two propagation environments: Indoor Hotspot (InH) and Urban Microcell (UMi). The normalized mean square error (NMSE), defined as $\mathrm{NMSE}=\mathbb{E}[||\hat{\mathbf{H}}-\mathbf{H}||_{F}^{2}/||\mathbf{H}||_{F}^{2}]$, is adopted as the evaluation metric. The 3D coordinates $(x,y,z)$ of the UE, BS, and RIS (in meters) are set to $(0,25,2)$, $(38,48,1)$, and $(40,50,2)$ for InH, and $(0,25,20)$, $(80,75,1)$, and $(70,85,10)$ for UMi.
For the channel configuration, the parameters are set as follows: $M=M_{h}\times M_{v}=8\times 8$, $N=N_{h}\times N_{v}=4\times 4$, and $K=17$. According to [18], the number of clusters follows a Poisson distribution, i.e., $C\sim\max\{\mathcal{P}(\lambda_{p}),1\}$, where $\lambda_{p}$ is set to $1.8$ for $f_{c}=28\ \text{GHz}$, and to $1.9$ for $f_{c}=73\ \text{GHz}$. The number of scatterers within each cluster follows a uniform distribution: $S_{c}\sim\mathcal{U}[1,30]$. $\zeta$ is set to $0.285$, and the angular spread is configured as $\sigma_{\varphi}=\sigma_{\phi}=5^{\circ}$. The environment-specific parameters used in (3) are configured as: InH(LOS): $n=1.73$, $\sigma_{x}=3.02$ dB, $b=0$; InH(NLOS): $n=3.19$, $\sigma_{x}=8.29$ dB, $b=0.06$; UMi(LOS): $n=1.98$, $\sigma_{x}=3.1$ dB, $b=0$; UMi(NLOS): $n=3.19$, $\sigma_{x}=8.2$ dB, $b=0$. The carrier frequency for all cases is $f_{0}=24.2$ GHz. Notably, for the InH scenario, the RIS is assumed to be sufficiently close to the BS such that the RIS–BS channel comprises only the LOS component, further highlighting the distinct characteristics between different scenarios. For the proposed method, the hyperparameters are set as follows: $T=1000$, the variance schedule $\{\beta_{t}\}$ increases linearly from $\beta_{0}=1\times 10^{-4}$ to $\beta_{T}=0.02$. The training set consists of 50,000 samples, while the validation and test sets contain 10,000 samples each.
@@ IV-B Results and Analysis
In the simulations, the proposed method is compared against the following baselines: (i) the classical least squares (LS) algorithm; (ii) the linear minimum mean square error (LMMSE) algorithm; (iii)DS_OMP [3], the double-structured orthogonal matching pursuit algorithm; (iv) CDRN [4], which employs a deep convolutional network to denoise LS estimation results; (v) DM-CNN1 [11], a method based on denoising diffusion probabilistic models (DDPMs) with a CNN-based network; and (vi) DM-CNN2, an enhanced version of DM-CNN1 with an increased number of convolutional layers and feature channels to ensure a fair comparison. For the proposed method, we implement three different network architectures: CNN2, U-Net, and the proposed BRCNet.
Fig. 3 illustrates the NMSE performance of different methods under varying SNR conditions. Specifically, in Fig. 3LABEL:sub@fig3:a at 0 dB, the proposed BRCNet-based method achieves performance gains of 14.70 dB, 6.12 dB, 8.69 dB, 2.03 dB, 3.91 dB, 1.65 dB, and 1.13 dB over LS, LMMSE, DS_OMP, CDRN, DM-CNN1, DM-CNN2, and Proposed-CNN2, respectively. Unlike CDRN’s direct-mapping method, the proposed method achieves improved performance through progressive denoising. Compared with DM-CNN1 and DM-CNN2, simulation results show that the proposed method is more effective. Moreover, the proposed BRCNet achieves performance highly comparable to U-Net, with the performance gap within 0.91 dB, while significantly reducing the number of parameters.
Table I compares the number of parameters and the number of multiply-accumulate operations (MACs) of DM-based methods. Although CNN1 has fewer parameters, its limited model capacity leads to poor performance in complex channel environments. In contrast, CNN2 has a similar parameter count to the BRCNet but still underperforms in terms of NMSE, highlighting the effectiveness of the proposed network architecture. Notably, BRCNet achieves performance comparable to U-Net with only 9.34% of its parameters, offering an excellent trade-off among model complexity, computational overhead, and estimation accuracy. Moreover, compared with CDRN, the proposed method can adapt to different SNRs without retraining, further reducing storage requirements and computational overhead in practical applications.
@@ V Conclusion
In this letter, we propose a DM-aided channel estimation method by formulating the estimation task as the reverse process of DMs, which adapts to varying noise levels. To reduce model complexity, we further design a parameter-efficient network, BRCNet, which achieves comparable performance with substantially fewer parameters. Simulation results demonstrate that the proposed method significantly outperforms baselines.

### v3

@@ III-C Network Architecture Design
1) Preprocess: To facilitate network processing, the complex-valued matrix $\mathbf{X}_{t}\in\mathbb{C}^{M\times K}$ is first decomposed into its real part $\Re(\mathbf{X}_{t})$ and imaginary part $\Im(\mathbf{X}_{t})$, and then reshaped into a three-dimensional tensor $\mathbf{X}^{\mathrm{real}}_{t}\in\mathbb{R}^{2\times M\times K}$.
2) $\text{Conv2d}(64,7\times 7)$: Convolutional layers have been widely adopted in image processing tasks due to their ability to capture local correlations between pixels. Since both channel matrices and images can be represented in matrix form, we apply a convolutional layer with 64 output channels and a $\times7$ kernel, denoted as $\text{Conv2d}(64,7\times 7)$, to effectively extract correlations among elements within the input.
3) Encoder: Motivated by the architectural efficiency and denoising effectiveness of DnCNN [Zhang2017], we adapt its design principles to channel estimation and introduce a novel BRC block. Specifically, this block consists of Batch normalization (BN) layers, ReLU activation functions, and Convolutional layers to extract multi-level features efficiently. To further enhance the model’s capability, we integrate the self-attention mechanism. This tailored design not only improves denoising performance but also reduces model complexity, making it well-suited for practical wireless communication scenarios.
Temporal information $t$ is embedded into each BRC block via a sinusoidal positional embedding vector $\mathbf{e}\in\mathbb{R}^{N_{t_{embed}}}$, enabling the model to capture temporal dependencies. The embedding vector is then projected to $\mathbf{e}_{project}\in\mathbb{R}^{N_{project}}$ to facilitate subsequent computations by a projection module consisting of a Swish activation function and a linear layer.
@@ IV-A Experiment Setup
[…]
Following the hyperparameter configuration in [Ho2020], we set the total diffusion steps to $T=1000$. The variance schedule $\{\beta_{t}\}$ increases linearly from $\beta_{0}=1\times 10^{-4}$ to $\beta_{T}=0.02$. Our dataset generation code is developed based on the open-source MATLAB simulator SimRIS [Basar2021]. For each scenario, the training set consists of 50,000 channel samples, while the validation and test sets each contain 10,000 channel samples. For training, the Adam optimizer is employed with a cosine warm-up learning rate schedule, where the initial learning rate is set to $1\times 10^{-4}$. The model is trained for 800 epochs with a batch size of 256.
@@ IV-B Results and Analysis
In simulations, the proposed method is compared with the following baselines: (i) the LMMSE estimator; (ii) DS_OMP [Wei2021]; (iii) CDRN [Liu2021]; (iv) LCDM(CNN1) [Fesl2024], which is based on denoising diffusion probabilistic models (DDPMs) with a low-complexity CNN-based backbone; (v) LCDM(CNN2) [Fesl2024], an enhanced version with an increased number of convolutional layers and feature channels to ensure a fair comparison; and (vi) GDM [Tong2024], a method based on guided DMs for RIS channel estimation. For the proposed method, we implement three network architectures: CNN2, the U-Net architecture in [Wu2024], and the proposed BRCNet.
Fig. 3 illustrates the NMSE performance under different SNR conditions. As a representative example, Fig. 3LABEL:sub@fig3:d (0 dB) is analyzed in detail. Unlike the direct-mapping approach of CDRN, the proposed method achieves superior performance through progressive denoising. Compared with LCDM(CNN1) and LCDM(CNN2), Proposed(BRCNet) achieves gains of 9.01 dB and 3.01 dB, respectively, demonstrating its superior effectiveness. By introducing the step-matching mechanism, our method better aligns with channel estimation and therefore outperforms GDM. Furthermore, the comparison between Proposed(CNN2) and Proposed(BRCNet) shows that the tailored BRCNet architecture is more effective than a simple CNN under the same parameter budget. Moreover, the proposed BRCNet achieves performance highly comparable to that of U-Net, with a performance gap of only 0.48 dB, while significantly reducing the number of parameters.
We further conduct an ablation study to evaluate the impact of different diffusion step numbers $T$ and the step-matching mechanism. By changing $T$ in the “Proposed(BRCNet)” configuration from 1000 to 200, we obtain the “T=200(BRCNet)” curve. The results show that $T=1000$ outperforms $T=200$ at low SNRs due to its finer noise-injection and denoising granularity, while the two settings yield similar performance at high SNRs, where only a few denoising steps are sufficient. We also remove the step-matching mechanism to obtain the “w/o step-matching” curve. Without step-matching, the sampling process starts from random Gaussian noise and does not utilize the received signal, resulting in SNR-insensitive performance. In contrast, with step-matching, sampling begins from the received signal itself, enabling effective denoising and yielding better SNR-dependent performance that better aligns with the requirements of the channel estimation task.
Table I compares the number of parameters and the number of multiply-accumulate operations (MACs) of different networks. Although CNN1 has fewer parameters, its limited capacity leads to inferior performance under complex channel conditions. CNN2 has a similar parameter count to that of BRCNet but still yields higher NMSE, highlighting the effectiveness of the proposed architecture. Notably, BRCNet achieves NMSE close to U-Net while using only 9.34% of its parameters, offering an excellent trade-off among model complexity, computational overhead, and estimation accuracy.
@@ V Conclusion
In this letter, we propose a DMs-aided channel estimation method by formulating the estimation task as the reverse process of DMs, which adapts to varying noise levels. To reduce model complexity, we further design a parameter-efficient network, BRCNet, which achieves comparable performance with substantially fewer parameters. Simulation results demonstrate that the proposed method significantly outperforms baselines.